Here is a word problem didn't know if i set it up right and solved it right
Three is added to a number and then the sum is multiplied by four, the result is 16
I got (n+3)4=16
n=3.25
Is this correct because it doesn't work out when you replace it with n

Answer by nerdybill(7384)

Your equation is correct:
(n+3)4=16
.
To solve for n, we start by dividing both sides by 4:
[(n+3)4]/4 = 16/4
(n+3) = 4
n+3 = 4
subtracting 3 from both sides:
n = 1